OpenStudy (anonymous):

Solve |4x-11| < 5. You must show all work. Use complete sentences to help me PLEASE!

OpenStudy (compassionate):

Hello, Davis. You'll be 16 soon, ya? Or has your birthday already passed? Anyways, to solve this, first solve like any normal inequality. 4x - 11 < 5 Okay, so just solve that for me, Davis.

OpenStudy (anonymous):

I'm 16 since the beginning of the year. And 4x<16 so x<4?

OpenStudy (anonymous):

My name is Ryan also. No need to call me Davis.

OpenStudy (compassionate):

Okay, so you would take 4x - 11 < 5, and since we know 4x - 11 are absolute values, and it is less than 5, we are going to add a minus five at the end. And that's really all there is too it! Then just solve it normally. -5 < 4x - 11 < 5

OpenStudy (anonymous):

how in the world did you get -5 < 4x - 11 < 5?

OpenStudy (jdoe0001):

one thing to keep in mind is that, when they say -> |4x-11| < 5 they really mean +1(4x-11) < 5 AND -1(4x-11) < 5 so, "x" takes those 2 values or as Compassionate already suggested, -5 < (4x-11) < 5 which is another way to write the same exact thing

OpenStudy (anonymous):

would you distribute this? +1(4x-11) < 5 AND -1(4x-11) < 5

OpenStudy (compassionate):

Basically you're solving for the original equation and it's inverse.

OpenStudy (compassionate):

This isn't making any sense to you, is it? :)

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Nope, thats why I'm here. Simply, I need work and a few sentences on this test review i need to turn in today... last minute too.

OpenStudy (anonymous):

there are like.... 2 more after this as well.

OpenStudy (jdoe0001):

you distribute it, and then solve the inequality, one will give you one value for "x", the other will give you a second value for "x", and "x" takes on both values

OpenStudy (anonymous):

hmm.... jdoe doesnt explain it as well as Compassionate does unfortunately.. lol sorry.

OpenStudy (jdoe0001):

hehhe, alright

OpenStudy (anonymous):

slope intercept is y-mx+b correct?

OpenStudy (jdoe0001):

y = mx+b, yes

OpenStudy (anonymous):

Change 3x + 2y = 10 to slope-intercept form. So it would be 2y= -3x+10?

OpenStudy (jdoe0001):

correct, and then divide both sides by "2"l, to isolate "y" on the left-hand-side

OpenStudy (compassionate):

No problem, Ryan You see, when we say things like |5|, what we're really saying is, "how many spaces from 0 is five?" Well, we know that 5 is positive five spaces and negative five spaces. So the absolute of five is simply positive and negative five on a graph. (see attachment 1 for example)[Screenshot 3] As, the absolute value if |-3| is, "how many spaces from 0 is -3? How many spaces from 0 is positive 3?" Well, it's positive and negative 3 spaces. Right? So back to your equation, it's asking: |4x-11| < 5 You have the absolute value: |4x-11| You're basically finding the positive and negative spaces. I'm being asked: "How many spaces to the right and left is 4x - 11 from 0?" So we know the positive spaces is 4x-11 And the negative spaces would be: -4x + 11 (because we multiplied by -1[or just take the inverse]) that makes sense, doesn't it? So now we're left with two inequalities. The positive one (or the original one) solves for the right side of the graph, and the negative one solves from the left side. So you have the two inequalities: 4x - 11 < 5 and -4x + 11 < 5 Now just solve each inequality and that's your solution!

OpenStudy (anonymous):

woah there...

OpenStudy (compassionate):

I hope this was helpful1

OpenStudy (jdoe0001):

2y= -3x+10 # divide both sides by 2 $$ 2y= -3x+10\\ \cfrac{2y}{2}= -\cfrac{3x}{2}+\cfrac{10}{2} $$

OpenStudy (anonymous):

so x<4 and x>3/2 is my final answer?

OpenStudy (anonymous):

and y= 3/2x+5?

OpenStudy (compassionate):

Ryan, you're a pro at this.

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

y= -3/2x+5 bear in mind the 3/2 is negative :)

OpenStudy (anonymous):

y-int is b in y+mx+b correct?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

double checking ;)

OpenStudy (anonymous):

y=mx+b Subtact 3x on both sides. 2y=-3x+10 Divide by 2. y= -3/2x +5 Slope: -3/2 y-int: 5

OpenStudy (jdoe0001):

 y-intercept, that is, where the function "touches" the y-axis

OpenStudy (anonymous):

Yes! Only 1 more question left on my review!

OpenStudy (anonymous):

i knew that part... im not that dumb. lol xD

OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1)?

OpenStudy (jdoe0001):

well, what would be the slope for that?

OpenStudy (anonymous):

use point-slope?

OpenStudy (jdoe0001):

\(m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (jdoe0001):

once you get the slope, first, you use the point-slope form then

OpenStudy (anonymous):

m is the slope. dunno how to find it. but -5/5 which is -1 for the rest.

OpenStudy (anonymous):

WAIT! Slope: m=-1! duh!

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

right?

OpenStudy (jdoe0001):

yes, now see this example http://www.coastal.edu/mathcenter/HelpPages/Line(Point-Slope)/img002.GIF

OpenStudy (jdoe0001):

you were given 2 points, use either on the point-slope form, then solve for "y"

OpenStudy (anonymous):

wut dafuq is dat....

OpenStudy (jdoe0001):

hehehe

OpenStudy (anonymous):

dammit, help?

OpenStudy (jdoe0001):

sorry I got really lagged for a bit :/